PartnerObserving Log for 2007 June 19-20

Observer: D. Thompson
LBC Support: A. DiPaola, R. Speziali, R. Bertram
Telescope Operator: D. Gonzalez

02:56 Dome open, problems with TCS startup? Enclosure not locked open?

03:03 Attempt to take bias images hung LBC (probably from TCS probs?). Doing lbckill/lbcstart.

03:23 Power up of HouseKeeping has failed several times with odd message. Cannot establish communication with the sensor (vacuum) - AdP looking into it.

04:22 AdP & DT went to PF to investigate a possible noise & check the system. While there, HK managed to power up OK. Rotators failed for a similar reason when powering up "other systems". AdP reccommends powering down and waiting ~20 minutes, it may be an overheating problem. Tout = 13.0 C.

04:27 Dome re-opened. All telescope systems up. Now we wait...

05:12 Re-powered HK - failed with same error: Cannot establish communication with the sensor (VACUUM)

06:22 LBC appears to have a hardware failure. AdP and RS will look at it tomorrow (Wednesday) in the daytime. We are dead for the night.

-- JohnHill - 22 Jun 2007
Topic revision: r3 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback